# Explainable Machine Learning Identification of Superconductivity from Single-Particle Spectral Functions

### v1

@@ Methods
Here, $\mathcal{H}$ denotes the Hopkins statistic [37], which measures the clustering tendency of the feature representation in the target domain; $\mathcal{M}$ denotes mutual information between the input and prediction in the target domain; $K$ is the number of classes for the normalization purpose. To obtain the Hopkins statistic, we define $\textbf{f}_{t}$ as the feature embeddings of all target domain samples, which is given by $\textbf{f}_{t}=[\mathrm{f}_{1},\mathrm{f}_{2},\ldots,\mathrm{f}_{n_{t}}]$, where $\mathrm{f}_{j}=G_{f}(\mathbf{x}_{j}^{t};\theta_{f})$. From $\textbf{f}_{t}$, we randomly sample $m=0.05n_{t}$ data points [38] without replacement to generate a set $R$. Additionally, we generate a set $U$ comprising $m$ data points sampled from a uniform distribution bounded by the minimum and maximum values along each feature dimension of $\textbf{f}_{t}$. We then compute two distance measures: $u_{k}$, the distance of samples in $U$ from their nearest neighbor in $R$, and $w_{k}$, the distance of samples in $R$ from their nearest neighbor in $R$. The Hopkins statistic is then defined as:
$\mathcal{H}=\frac{\sum^{m}_{k=1}u_{k}}{\sum^{m}_{k=1}u_{k}+\sum^{m}_{k=1}w_{k}}$ (7)
The mutual information is defined as:
@@
where $H(\cdot)$ denotes the information entropy. The optimized model is then trained using the labeled simulated training set and all unlabeled experimental spectra. The model checkpoint with the best TS is collected for test set predictions.
.5 Occlusion-based attribution
By moving the occluding patch along the energy axis, we quantify the saliency for each spectrum sample at energy level $\omega=\nu$ as:
where the occluded spectrum of the original spectrum $A(\mathbf{k},\omega)$, with respect to a single baseline spectrum ($A_{0}$) at energy $\nu$, is defined as
Here, $\bar{p}_{\textrm{SC}}[\cdot]$ represents the ML model mapping from an input ARPES spectrum to the predicted SC probability $\bar{p}_{\textrm{SC}}$. Since we focus on features with positive contributions, a ReLU function is applied to filter out the distraction of negative values. Following recent feature attribution studies [39, 40], we choose various occlusion baselines to ensure robust explanation performance across different pixel intensities (see details in Supplementary Note 6).
@@ Acknowledgements
We thank Mingda Li for insightful discussions. X.C., Y.S., and Y.W. acknowledge support from the Air Force Office of Scientific Research Young Investigator Program under grant FA9550-23-1-0153. Y.H. acknowledges support from NSF CAREER award No. DMR-2239171. J.Y. was partially supported by a seed fund from the Yale Provost’s Office. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award BES-ERCAP0023181.

### v2

@@ IV.4 Model evaluation and hyperparameter tuning
Here, $\mathcal{H}$ denotes the Hopkins statistic [44], which measures the clustering tendency of the feature representation in the target domain; $\mathcal{M}$ denotes mutual information between the input and prediction in the target domain; $K$ is the number of classes for the normalization purpose. To obtain the Hopkins statistic, we define $\textbf{f}_{t}$ as the feature embeddings of all target domain samples, which is given by $\textbf{f}_{t}=[\mathrm{f}_{1},\mathrm{f}_{2},\ldots,\mathrm{f}_{n_{t}}]$, where $\mathrm{f}_{j}=G_{f}(\mathbf{x}_{j}^{t};\theta_{f})$. From $\textbf{f}_{t}$, we randomly sample $m=0.05n_{t}$ data points [45], without replacement to generate a set $R$. Additionally, we generate a set $U$ comprising $m$ data points sampled from a uniform distribution bounded by the minimum and maximum values along each feature dimension of $\textbf{f}_{t}$. We then compute two distance measures: $u_{k}$, the distance of samples in $U$ from their nearest neighbor in $R$, and $w_{k}$, the distance of samples in $R$ from their nearest neighbor in $R$. The Hopkins statistic is then defined as:
$\mathcal{H}=\frac{\sum^{m}_{k=1}u_{k}}{\sum^{m}_{k=1}u_{k}+\sum^{m}_{k=1}w_{k}}\,.$ (12)
The mutual information is used to discern both the prediction confidence and diversity, as described in Yang et al. [43], and is defined as follows:
@@
where $H(\cdot)$ denotes the information entropy. With all the hyperparameter optimized, the model is then trained using all the labeled simulated training set and unlabeled experimental spectra. The model checkpoint with the best TS is collected for test set predictions.
Data Availability
All the ARPES data that support the findings of this study are deposited in the Figshare repository (https://doi.org/10.6084/m9.figshare.25439632.v1).
Code Availability
The training script used to produce the findings of this study is publicly available via GitHub at https://github.com/Liu-group/ARPES and via Figshare at https://doi.org/10.6084/m9.figshare.25439623.v1.
@@ Acknowledgements
We thank Mingda Li for insightful discussions. X.C. and F.L. acknowledge support by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Early Career Award No. DE-SC0025345. X.C., Y.S., and Y.W. acknowledge support from the Air Force Office of Scientific Research Young Investigator Program under grant FA9550-23-1-0153. Y.H. acknowledges support from NSF CAREER award No. DMR-2239171. J.Y. was partially supported by a seed fund from the Yale Provost’s Office. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award BES-ERCAP0031226.